Which number completes this sequence? 123, 117, 108, 99,
Answer by Edwin McCravy(20056)

123,  117, 108, 99, ?, ?, ?, ?, ?, ?, ?, ?, ? 

Start with 123: 

123, find the sum of the digits 1+2+3=6, Subtract 123-6 = 117

117, find the sum of the digits 1+1+7=9, Subtract 117-9 = 108

108, find the sum of the digits 1+0+8=9, Subtract 108-9 = 99

99, find the sum of the digits 9+9=18, Subtract 99-18 = 81

81, find the sum of the digits 8+1=9, Subtract 81-9 = 72

72, find the sum of the digits 7+2=9, Subtract 72-9 = 63

63, find the sum of the digits 6+3=9, Subtract 63-9 = 54

54, find the sum of the digits 5+4=9, Subtract 54-9 = 45

45, find the sum of the digits 4+5=9, Subtract 45-9 = 36

36, find the sum of the digits 3+6=9, Subtract 36-9 = 27

27, find the sum of the digits 2+7=9, Subtract 27-9 = 18

18, find the sum of the digits 1+8=9, Subtract 18-9 =  9

9, find the sum of the digits 9, Subtract 9-9 =  0

Edwin